# Koby Hayashi

### v1

@@ 4.3.1 Leverage Score Sampling for NLS Problems
$={\bm{\mathbf{{z}}}}^{\sf T}\Big{[}{\bm{\mathbf{{U}}}}_{A}^{\sf T}({\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}_{nls}-{\bm{\mathbf{{b}}}})-{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}({\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}_{nls}-{\bm{\mathbf{{b}}}})\Big{]}={\bm{\mathbf{{z}}}}^{\sf T}\Big{[}{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{r}}}}_{nls}-{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}{\bm{\mathbf{{r}}}}_{nls}\Big{]}.$
Recall that ${\bm{\mathbf{{r}}}}_{nls}={\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}_{nls}-{\bm{\mathbf{{b}}}}$ is the NLS residual vector. Combining the previous two equations back into Lemma 3 we have
${\bm{\mathbf{{z}}}}^{\sf T}\Big{[}{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{r}}}}_{nls}-{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}{\bm{\mathbf{{r}}}}_{nls}\Big{]}\geq\sigma_{min}({\bm{\mathbf{{S}}}}{\bm{\mathbf{{U}}}}_{A})\|{\bm{\mathbf{{z}}}}\|_{2}^{2}$
$\rightarrow\|{\bm{\mathbf{{z}}}}\|_{2}\Big{\|}{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{r}}}}_{nls}-{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}{\bm{\mathbf{{r}}}}_{nls}\Big{\|}_{2}\geq\sigma_{min}({\bm{\mathbf{{S}}}}{\bm{\mathbf{{U}}}}_{A})\|{\bm{\mathbf{{\Sigma}}}}_{A}{\bm{\mathbf{{V}}}}_{A}(\hat{{\bm{\mathbf{{x}}}}}_{nls}-{\bm{\mathbf{{x}}}}_{nls})\|_{2}^{2}.$
@@
so that both SC1, as in Theorem 7 with $\epsilon_{s}=1-1/\sqrt{2}$ meaning that $\sigma_{min}({\bm{\mathbf{{S}}}}{\bm{\mathbf{{U}}}}_{A})\geq 1/\sqrt{2}$, and SC2 as in Equation 29 as in Lemma 9 with ${\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{r}}}}_{cls}$ being the matrix product to approximate. Each SC holds with at least probability $21-\delta/$ 2 so they both hold with at least probability $1-\delta$. Picking up from where we left off and assuming that both SC’s hold we have
$\sigma_{min}({\bm{\mathbf{{S}}}}{\bm{\mathbf{{U}}}}_{A})\|{\bm{\mathbf{{\Sigma}}}}_{A}{\bm{\mathbf{{V}}}}_{A}(\hat{{\bm{\mathbf{{x}}}}}_{nls}-{\bm{\mathbf{{x}}}}_{nls})\|_{2}^{2}\leq\|{\bm{\mathbf{{z}}}}\|_{2}\Big{\|}{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{r}}}}_{nls}-{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}{\bm{\mathbf{{r}}}}_{nls}\Big{\|}_{2}$
$\bm{\rightarrow_{SC^{\prime}s}}\frac{1}{\sqrt{2}}\sigma_{min}({\bm{\mathbf{{A}}}})\|(\hat{{\bm{\mathbf{{x}}}}}_{nls}-{\bm{\mathbf{{x}}}}_{nls})\|_{2}\leq\frac{\sqrt{\epsilon_{r}}\|{\bm{\mathbf{{r}}}}_{nls}\|_{2}}{\sqrt{2}}\rightarrow\|\hat{{\bm{\mathbf{{x}}}}}_{nls}-{\bm{\mathbf{{x}}}}_{nls}\|_{2}\leq\frac{\sqrt{\epsilon_{r}}\|{\bm{\mathbf{{r}}}}_{nls}\|_{2}}{\sigma_{min}({\bm{\mathbf{{A}}}})}.$
Where the notation $\bm{\rightarrow_{SC^{\prime}s}}$ is used to emphasize the usage of the two SC’s. This concludes the proof of Theorem 2. ∎
@@ 4.3.2 Analysis of Hybrid Sampling
This section presents our theoretical results for Hybrid Sampling. According to the proof of Theorem 2 to show that Hybrid Sampling works for OLS and NLS problems we need only show that the Hybrid Sampling matrix satisfies the two SC’s in Equation 28 and Equation 29. Applying these results in the proof structure used for Theorem 2 will yield sampling complexities, theoretical guarantees, and algorithms for Hybrid Sampling and solving NLS problems.
The first Structural Condition for Hybrid Sampling is given by Lemma 4.
Lemma 4.
Given ${\bm{\mathbf{{A}}}}\in\mathbb{R}^{m\times k}$ consider its SVD ${\bm{\mathbf{{U}}}}_{A}{\bm{\mathbf{{\Sigma}}}}_{A}{\bm{\mathbf{{V}}}}_{A}^{\sf T}$ and its row leverage scores $l_{i}({\bm{\mathbf{{A}}}})$ for each row $i\in[m]$, where $[m]$ denotes the set of integers from 1 to $m$. Let $0<\tau\leq 1$ be the deterministic sampling parameter such that if $\tau\leq l_{i}({\bm{\mathbf{{A}}}})/k$ then $i\in\mathscr{I}_{D}$ and $\mathscr{I}_{R}=[m]\setminus\mathscr{I}_{D}$. Let ${\bm{\mathbf{{S}}}}_{H}\in\mathbb{R}^{s\times m}$ be a row sampling and rescaling matrix contructed via sampling with replacement on $\mathscr{I}_{R}$ with $s_{R}$ samples drawn according to the renormalized leverage scores $l_{i}({\bm{\mathbf{{A}}}})$, $s_{D}$ deterministic samples taken from $\mathscr{I}_{D}$ where $|\mathscr{I}_{D}|=s_{D}$, and $s=s_{D}+s_{R}$ Define $\theta=\sum_{i\in\mathscr{I}_{D}}l_{i}({\bm{\mathbf{{A}}}})$ and $\xi=k-\theta$. If $\frac{144\xi\log(2k/\delta)}{\epsilon_{s}^{2}}<s_{R}$ then the following equation holds with at least probability $1-\delta$
@@
This proof is a modification of the proofs in [45, 31] for proving the analogous statement for standard leverage score sampling, which apply a Matrix Chernoff Bound. A statement of the Matrix Chernoff Bound is available in the Appendix as Theorem 10. This theorem allows us to reason about the quantity $Pr\big{[}\|{\bm{\mathbf{{W}}}}\|_{2}>\epsilon\big{]}$ where ${\bm{\mathbf{{W}}}}=\frac{1}{s_{R}}\sum_{j=1}^{s_{R}}{\bm{\mathbf{{X}}}}_{j}$ and ${\bm{\mathbf{{X}}}}_{j}$ are draws of an independent, symmetric random matrix ${\bm{\mathbf{{X}}}}\in\mathbb{R}^{k\times k}$. Additionally let $\mathscr{S}_{R}$ be the set of samples or draws of ${\bm{\mathbf{{X}}}}$ such that $|\mathscr{S}_{R}|=s_{R}$. To prove lemma 4 we set ${\bm{\mathbf{{W}}}}={\bm{\mathbf{{I}}}}-{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{S}}}}_{H}^{\sf T}{\bm{\mathbf{{S}}}}_{H}{\bm{\mathbf{{U}}}}_{A}$, where ${\bm{\mathbf{{S}}}}_{H}$ is given by Equation 25.
Proof.
@@
We assume an appropriate row permutation conformal with eq. 25 such that ${\bm{\mathbf{{U}}}}_{A}=\begin{bmatrix}{\bm{\mathbf{{U}}}}_{D}\\ {\bm{\mathbf{{U}}}}_{R}\end{bmatrix}$ where ${\bm{\mathbf{{U}}}}_{D}\in\mathbb{R}^{(s_{D}\times k)}$ and ${\bm{\mathbf{{U}}}}_{R}\in\mathbb{R}^{(m-s_{D})\times k}$. Define ${\bm{\mathbf{{G}}}}_{D}=({\bm{\mathbf{{S}}}}_{D}{\bm{\mathbf{{U}}}}_{D})^{\sf T}{\bm{\mathbf{{S}}}}_{D}{\bm{\mathbf{{U}}}}_{D}={\bm{\mathbf{{U}}}}_{D}^{\sf T}{\bm{\mathbf{{U}}}}_{D}$ and ${\bm{\mathbf{{G}}}}_{R}={\bm{\mathbf{{U}}}}_{R}^{\sf T}{\bm{\mathbf{{U}}}}_{R}$. Then we have $({\bm{\mathbf{{S}}}}_{H}{\bm{\mathbf{{U}}}}_{A})^{\sf T}({\bm{\mathbf{{S}}}}_{H}{\bm{\mathbf{{U}}}}_{A})={\bm{\mathbf{{G}}}}_{D}+({\bm{\mathbf{{S}}}}_{R}{\bm{\mathbf{{U}}}}_{R})^{\sf T}({\bm{\mathbf{{S}}}}_{R}{\bm{\mathbf{{U}}}}_{R})$. Let ${\bm{\mathbf{{X}}}}_{j}={\bm{\mathbf{{I}}}}_{k}-{\bm{\mathbf{{G}}}}_{D}-\frac{1}{\tilde{\pi}_{j}}{\bm{\mathbf{{q}}}}_{j}{\bm{\mathbf{{q}}}}_{j}^{\sf T}$ be a sequence of draws from ${\bm{\mathbf{{X}}}}={\bm{\mathbf{{I}}}}_{k}-{\bm{\mathbf{{G}}}}_{D}-\frac{1}{\tilde{\pi}}{\bm{\mathbf{{q}}}}{\bm{\mathbf{{q}}}}^{\sf T}$ for $j\in[s_{R}]$. This implies $\tilde{\pi}_{j}=\tilde{p}_{i}$ and ${\bm{\mathbf{{q}}}}_{j}={\bm{\mathbf{{u}}}}_{i}$ if the $i$th row of ${\bm{\mathbf{{U}}}}$ is the $j$th sample, drawn according to the distribution given by $\tilde{p}_{i}=\|{\bm{\mathbf{{u}}}}_{i}\|_{2}^{2}/\|{\bm{\mathbf{{U}}}}_{R}\|_{F}^{2}=\|{\bm{\mathbf{{u}}}}_{i}\|_{2}^{2}/\xi$, the corresponding leverage score probability for $i\in\mathscr{I}_{R}$. Note that for any $i$ we have $\frac{\|{\bm{\mathbf{{u}}}}_{i}\|_{2}^{2}}{\tilde{p}_{i}}=\frac{\|{\bm{\mathbf{{u}}}}_{i}\|_{2}^{2}}{\|{\bm{\mathbf{{u}}}}_{i}\|_{2}^{2}}\xi=\xi$.
Then to obtain ${\bm{\mathbf{{W}}}}={\bm{\mathbf{{I}}}}_{k}-({\bm{\mathbf{{S}}}}_{H}{\bm{\mathbf{{U}}}}_{A})^{\sf T}{\bm{\mathbf{{S}}}}_{H}{\bm{\mathbf{{U}}}}_{A}$ as a sum of ${\bm{\mathbf{{X}}}}_{j}$’s, we observe
$\frac{1}{s_{R}}\sum_{j=1}^{s_{R}}{\bm{\mathbf{{X}}}}_{j}=\frac{1}{s_{R}}\sum_{j=1}^{s_{R}}\Big{(}{\bm{\mathbf{{I}}}}_{k}-{\bm{\mathbf{{G}}}}_{D}-\frac{1}{\tilde{\pi}_{j}}{\bm{\mathbf{{q}}}}_{j}{\bm{\mathbf{{q}}}}_{j}^{\sf T}\Big{)}={\bm{\mathbf{{I}}}}_{k}-{\bm{\mathbf{{G}}}}_{D}-\sum_{j=1}^{s_{R}}\frac{{\bm{\mathbf{{q}}}}_{j}}{\sqrt{s_{R}\tilde{\pi}_{j}}}\frac{{\bm{\mathbf{{q}}}}_{j}^{\sf T}}{\sqrt{s_{R}\tilde{\pi}_{j}}}$
$={\bm{\mathbf{{I}}}}_{k}-{\bm{\mathbf{{G}}}}_{D}-({\bm{\mathbf{{S}}}}_{R}{\bm{\mathbf{{U}}}}_{R})^{\sf T}{\bm{\mathbf{{S}}}}_{R}{\bm{\mathbf{{U}}}}_{R}={\bm{\mathbf{{I}}}}_{k}-({\bm{\mathbf{{S}}}}_{H}{\bm{\mathbf{{U}}}}_{A})^{\sf T}{\bm{\mathbf{{S}}}}_{H}{\bm{\mathbf{{U}}}}_{A}={\bm{\mathbf{{W}}}}.$
@@
So ${\bm{\mathbf{{I}}}}_{k}-({\bm{\mathbf{{S}}}}_{H}{\bm{\mathbf{{U}}}}_{A})^{\sf T}{\bm{\mathbf{{S}}}}_{H}{\bm{\mathbf{{U}}}}_{A}$ is a sum of symmetric, independent random matrices. Next, to apply Theorem 10 we need to verify three conditions: $\mathbb{E}[{\bm{\mathbf{{X}}}}]=0$, $\|{\bm{\mathbf{{X}}}}\|_{2}\leq\gamma$, and $\|\mathbb{E}[{\bm{\mathbf{{X}}}}^{\sf T}{\bm{\mathbf{{X}}}}]\|_{2}\leq\nu^{2}$, where $\gamma$ and $\nu^{2}$ are bounds used in Theorem 10 to be derived.
@@
$\mathbb{E}[{\bm{\mathbf{{X}}}}]=\mathbb{E}[{\bm{\mathbf{{I}}}}_{k}-({\bm{\mathbf{{S}}}}_{D}{\bm{\mathbf{{U}}}}_{D})^{\sf T}({\bm{\mathbf{{S}}}}_{D}{\bm{\mathbf{{U}}}}_{D})-\frac{1}{\tilde{p}}{\bm{\mathbf{{q}}}}{\bm{\mathbf{{q}}}}^{\sf T}]={\bm{\mathbf{{G}}}}_{R}-\mathbb{E}[\frac{1}{\tilde{\pi}}{\bm{\mathbf{{q}}}}{\bm{\mathbf{{q}}}}^{\sf T}]={\bm{\mathbf{{0}}}}_{k\times k}$
Second, bound $\|{\bm{\mathbf{{X}}}}\|_{2}\leq\gamma$. Observe $\frac{1}{\tilde{\pi}}{\bm{\mathbf{{q}}}}{\bm{\mathbf{{q}}}}^{\sf T}$ is a rank-1 matrix with its 2-norm bounded by $\max_{i\in{\cal I}_{R}}\frac{\|{\bm{\mathbf{{u}}}}_{i}\|_{2}^{2}}{\tilde{p}_{i}}=\xi$, then
$\|{\bm{\mathbf{{X}}}}\|_{2}\leq\|{\bm{\mathbf{{I}}}}_{k}-({\bm{\mathbf{{S}}}}_{D}{\bm{\mathbf{{U}}}}_{D})^{\sf T}({\bm{\mathbf{{S}}}}_{D}{\bm{\mathbf{{U}}}}_{D})\|_{2}+\|\frac{1}{\tilde{p}}{\bm{\mathbf{{q}}}}{\bm{\mathbf{{q}}}}^{\sf T}\|_{2}\leq 1+\xi.$
@@
Finally, bound $\|\mathbb{E}[{\bm{\mathbf{{X}}}}^{\sf T}{\bm{\mathbf{{X}}}}]\|\leq\nu^{2}$. By expanding $\mathbb{E}[{\bm{\mathbf{{X}}}}^{\sf T}{\bm{\mathbf{{X}}}}]$ and using the fact that
@@
$Pr[\|{\bm{\mathbf{{I}}}}-{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{S}}}}_{H}^{\sf T}{\bm{\mathbf{{S}}}}_{H}{\bm{\mathbf{{U}}}}_{A}\|_{2}^{2}\geq\epsilon_{s}]=\delta\leq 2k\exp\Big{[}-s\epsilon_{s}^{2}/\Big{(}2(|\xi-1|)+2(1+\xi)\epsilon_{s}/3\Big{)}\Big{]}$ (35)
for some desired probability $\delta$ and desired error tolerance $\epsilon_{s}$. By choosing $s_{R}\geq\frac{144\xi\log(2k/\delta)}{\epsilon_{s}^{2}}$ we have with at least probability $1-\delta$

### v2

@@ 4.3.1 Leverage Score Sampling for NLS Problems
$={\bm{\mathbf{{z}}}}^{\sf T}\Big{[}{\bm{\mathbf{{U}}}}_{A}^{\sf T}({\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}_{nls}-{\bm{\mathbf{{b}}}})-{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}({\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}_{nls}-{\bm{\mathbf{{b}}}})\Big{]}$
$={\bm{\mathbf{{z}}}}^{\sf T}\Big{[}{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{r}}}}_{nls}-{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}{\bm{\mathbf{{r}}}}_{nls}\Big{]}.$
Recall that ${\bm{\mathbf{{r}}}}_{nls}={\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}_{nls}-{\bm{\mathbf{{b}}}}$ is the NLS residual vector. Combining the previous two equations back into Lemma 4.1 we have
${\bm{\mathbf{{z}}}}^{\sf T}\Big{[}{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{r}}}}_{nls}-{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}{\bm{\mathbf{{r}}}}_{nls}\Big{]}$ $\geq\sigma_{min}({\bm{\mathbf{{S}}}}{\bm{\mathbf{{U}}}}_{A})^{2}\|{\bm{\mathbf{{z}}}}\|_{2}^{2},$
which implies
(29) $\|{\bm{\mathbf{{z}}}}\|_{2}\Big{\|}{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{r}}}}_{nls}-{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{S}}}}^{\sf T}{\bm{\mathbf{{S}}}}{\bm{\mathbf{{r}}}}_{nls}\Big{\|}_{2}\geq\sigma_{min}({\bm{\mathbf{{S}}}}{\bm{\mathbf{{U}}}}_{A})^{2}\|{\bm{\mathbf{{\Sigma}}}}_{A}{\bm{\mathbf{{V}}}}_{A}(\hat{{\bm{\mathbf{{x}}}}}_{nls}-{\bm{\mathbf{{x}}}}_{nls})\|_{2}^{2}.$
@@
so that both SC1, as in Theorem F.1 with $\epsilon_{s}=1-1/\sqrt{2}$ meaning that $\sigma_{min}({\bm{\mathbf{{S}}}}{\bm{\mathbf{{U}}}}_{A})^{2}\geq 1/\sqrt{2}$, and SC2, as in Equation 23 as in Lemma F.3 with ${\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{r}}}}_{nls}$ being the matrix product to approximate, both hold with high probability. With these parameters each SC holds with at least probability $21-\delta/$ 2 so they both hold with at least probability $1-\delta$. Thus, from Equation 29 we have that if both SC’s hold, then
$\frac{1}{\sqrt{2}}\sigma_{min}({\bm{\mathbf{{A}}}})\|(\hat{{\bm{\mathbf{{x}}}}}_{nls}-{\bm{\mathbf{{x}}}}_{nls})\|_{2}\leq\frac{\sqrt{\epsilon_{r}}\|{\bm{\mathbf{{r}}}}_{nls}\|_{2}}{\sqrt{2}},$
from which the claimed upper bound immediately follows.
Discussion
As stated previously Theorem 2.1 is applicable to all problems of the form
(30) $\min_{{\bm{\mathbf{{x}}}}\in\mathscr{C}}\|{\bm{\mathbf{{A}}}}{\bm{\mathbf{{x}}}}-{\bm{\mathbf{{b}}}}\|_{2}.$
with overdetermined, full rank ${\bm{\mathbf{{A}}}}$, i.e., all convex LS problems. This is due to the fact that Eq. 28 is a result of convexity. Additionally, there are other types of sketches which can be used. For example, the Subsampled Randomized Hadamard Transform discussed in [14] and the ‘sparse embedding matrices’ from [56] can be applied. This is because these sketches can satisfy the two SC’s with an appropriate number of samples.
@@ 4.3.2 Analysis of Hybrid Sampling
This section presents our theoretical results for Hybrid Sampling. According to the proof of Theorem 2.1 to show that Hybrid Sampling works for OLS and NLS problems we need only show that the Hybrid Sampling matrix satisfies the two SC’s in Equation 22 and Equation 23. Applying these results in the proof structure used for Theorem 2.1 will yield sampling complexities, theoretical guarantees, and algorithms for Hybrid Sampling and solving NLS problems.
The first Structural Condition for Hybrid Sampling is given by Lemma 4.3.
Lemma 4.3.
Given ${\bm{\mathbf{{A}}}}\in\mathbb{R}^{m\times k}$ consider its thin SVD ${\bm{\mathbf{{U}}}}_{A}{\bm{\mathbf{{\Sigma}}}}_{A}{\bm{\mathbf{{V}}}}_{A}^{\sf T}$ and its row leverage scores $l_{i}({\bm{\mathbf{{A}}}})$ for each row $i\in[m]$, where $[m]$ denotes the set of integers from 1 to $m$. Let $\mathscr{I}_{D}$ be the set of deterministically included rows and define $\mathscr{I}_{R}=[m]\setminus\mathscr{I}_{D}$ as the set of rows to be sampled from. Let ${\bm{\mathbf{{S}}}}_{H}\in\mathbb{R}^{s\times m}$ be a row sampling and rescaling matrix contructed via sampling with replacement on $\mathscr{I}_{R}$ with $s_{R}$ samples drawn according to the renormalized leverage scores $l_{i}({\bm{\mathbf{{A}}}})$, $s_{D}$ deterministic samples taken from $\mathscr{I}_{D}$ where $|\mathscr{I}_{D}|=s_{D}$, and $s=s_{D}+s_{R}$ Define $\theta=\sum_{i\in\mathscr{I}_{D}}l_{i}({\bm{\mathbf{{A}}}})$ and $\xi=k-\theta$. If $\frac{144\xi\log(2k/\delta)}{\epsilon_{s}^{2}}<s_{R}$ then the following equation holds with probability at least $1-\delta$
@@
This proof is a modification of the proofs in [56, 41] for proving the analogous statement for standard leverage score sampling, which apply a Matrix Chernoff Bound. A statement of the Matrix Chernoff Bound is available in the Appendix as Theorem F.5. This theorem allows us to reason about the quantity $Pr\big{[}\|{\bm{\mathbf{{W}}}}\|_{2}>\epsilon\big{]}$ where ${\bm{\mathbf{{W}}}}=\frac{1}{s_{R}}\sum_{j=1}^{s_{R}}{\bm{\mathbf{{X}}}}_{j}$ and ${\bm{\mathbf{{X}}}}_{j}$ are draws of an independent, symmetric random matrix ${\bm{\mathbf{{X}}}}\in\mathbb{R}^{k\times k}$. Additionally let $\mathscr{S}_{R}$ be the set of samples or draws of ${\bm{\mathbf{{X}}}}$ such that $|\mathscr{S}_{R}|=s_{R}$. To prove Lemma 4.3 we set ${\bm{\mathbf{{W}}}}={\bm{\mathbf{{I}}}}-{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{S}}}}_{H}^{\sf T}{\bm{\mathbf{{S}}}}_{H}{\bm{\mathbf{{U}}}}_{A}$, where ${\bm{\mathbf{{S}}}}_{H}$ is given by Eq. 20.
Proof 4.4.
@@
We assume an appropriate row permutation conformal with Eq. 20 such that ${\bm{\mathbf{{U}}}}_{A}=\begin{bmatrix}{\bm{\mathbf{{U}}}}_{D}\\ {\bm{\mathbf{{U}}}}_{R}\end{bmatrix}$ where ${\bm{\mathbf{{U}}}}_{D}\in\mathbb{R}^{(s_{D}\times k)}$ and ${\bm{\mathbf{{U}}}}_{R}\in\mathbb{R}^{(m-s_{D})\times k}$. Define ${\bm{\mathbf{{G}}}}_{D}=({\bm{\mathbf{{S}}}}_{D}{\bm{\mathbf{{U}}}}_{D})^{\sf T}{\bm{\mathbf{{S}}}}_{D}{\bm{\mathbf{{U}}}}_{D}={\bm{\mathbf{{U}}}}_{D}^{\sf T}{\bm{\mathbf{{U}}}}_{D}$ and ${\bm{\mathbf{{G}}}}_{R}={\bm{\mathbf{{U}}}}_{R}^{\sf T}{\bm{\mathbf{{U}}}}_{R}$. Then we have $({\bm{\mathbf{{S}}}}_{H}{\bm{\mathbf{{U}}}}_{A})^{\sf T}({\bm{\mathbf{{S}}}}_{H}{\bm{\mathbf{{U}}}}_{A})={\bm{\mathbf{{G}}}}_{D}+({\bm{\mathbf{{S}}}}_{R}{\bm{\mathbf{{U}}}}_{R})^{\sf T}({\bm{\mathbf{{S}}}}_{R}{\bm{\mathbf{{U}}}}_{R})$. Let ${\bm{\mathbf{{q}}}}$ be a random row sample drawn from ${\bm{\mathbf{{U}}}}_{R}$ according to the distribution given by $\tilde{p}_{i}=\|{\bm{\mathbf{{u}}}}_{i}\|_{2}^{2}/\|{\bm{\mathbf{{U}}}}_{R}\|_{F}^{2}=\|{\bm{\mathbf{{u}}}}_{i}\|_{2}^{2}/\xi$, the corresponding leverage score probability for $i\in\mathscr{I}_{R}$, and let $\tilde{\pi}$ be ${\bm{\mathbf{{q}}}}$’s leverage score. Define ${\bm{\mathbf{{q}}}}_{j}$ for $j\in[s_{R}]$ to be a sequence of independent draws of ${\bm{\mathbf{{q}}}}$, so that ${\bm{\mathbf{{q}}}}_{j}={\bm{\mathbf{{u}}}}_{i}$ and $\tilde{\pi}_{j}=\tilde{p}_{i}$ if the $i$th row of ${\bm{\mathbf{{U}}}}$ is the $j$th sample. Let ${\bm{\mathbf{{X}}}}_{j}={\bm{\mathbf{{I}}}}_{k}-{\bm{\mathbf{{G}}}}_{D}-\frac{1}{\tilde{\pi}_{j}}{\bm{\mathbf{{q}}}}_{j}{\bm{\mathbf{{q}}}}_{j}^{\sf T}$, which is an independent, symmetric random matrix drawn from ${\bm{\mathbf{{X}}}}={\bm{\mathbf{{I}}}}_{k}-{\bm{\mathbf{{G}}}}_{D}-\frac{1}{\tilde{\pi}}{\bm{\mathbf{{q}}}}{\bm{\mathbf{{q}}}}^{\sf T}$. Note that for any $i$ we have $\frac{\|{\bm{\mathbf{{u}}}}_{i}\|_{2}^{2}}{\tilde{p}_{i}}=\frac{\|{\bm{\mathbf{{u}}}}_{i}\|_{2}^{2}}{\|{\bm{\mathbf{{u}}}}_{i}\|_{2}^{2}}\xi=\xi$.
Then
${\bm{\mathbf{{W}}}}=\frac{1}{s_{R}}\sum_{j=1}^{s_{R}}{\bm{\mathbf{{X}}}}_{j}=\frac{1}{s_{R}}\sum_{j=1}^{s_{R}}\Big{(}{\bm{\mathbf{{I}}}}_{k}-{\bm{\mathbf{{G}}}}_{D}-\frac{1}{\tilde{\pi}_{j}}{\bm{\mathbf{{q}}}}_{j}{\bm{\mathbf{{q}}}}_{j}^{\sf T}\Big{)}={\bm{\mathbf{{I}}}}_{k}-{\bm{\mathbf{{G}}}}_{D}-\sum_{j=1}^{s_{R}}\frac{{\bm{\mathbf{{q}}}}_{j}}{\sqrt{s_{R}\tilde{\pi}_{j}}}\frac{{\bm{\mathbf{{q}}}}_{j}^{\sf T}}{\sqrt{s_{R}\tilde{\pi}_{j}}}$
$={\bm{\mathbf{{I}}}}_{k}-{\bm{\mathbf{{G}}}}_{D}-({\bm{\mathbf{{S}}}}_{R}{\bm{\mathbf{{U}}}}_{R})^{\sf T}{\bm{\mathbf{{S}}}}_{R}{\bm{\mathbf{{U}}}}_{R}={\bm{\mathbf{{I}}}}_{k}-({\bm{\mathbf{{S}}}}_{H}{\bm{\mathbf{{U}}}}_{A})^{\sf T}{\bm{\mathbf{{S}}}}_{H}{\bm{\mathbf{{U}}}}_{A}.$
@@
So ${\bm{\mathbf{{I}}}}_{k}-({\bm{\mathbf{{S}}}}_{H}{\bm{\mathbf{{U}}}}_{A})^{\sf T}{\bm{\mathbf{{S}}}}_{H}{\bm{\mathbf{{U}}}}_{A}$ is a sum of symmetric, independent random matrices. Next, to apply Theorem F.5 we need to verify three conditions: $\mathbb{E}[{\bm{\mathbf{{X}}}}]=0$, $\|{\bm{\mathbf{{X}}}}\|_{2}\leq\gamma$, and $\|\mathbb{E}[{\bm{\mathbf{{X}}}}^{\sf T}{\bm{\mathbf{{X}}}}]\|_{2}\leq\nu^{2}$, where $\gamma$ and $\nu^{2}$ are bounds used in Theorem F.5 to be derived.
@@
$\mathbb{E}[{\bm{\mathbf{{X}}}}]=\mathbb{E}[{\bm{\mathbf{{I}}}}_{k}-({\bm{\mathbf{{S}}}}_{D}{\bm{\mathbf{{U}}}}_{D})^{\sf T}({\bm{\mathbf{{S}}}}_{D}{\bm{\mathbf{{U}}}}_{D})-\frac{1}{\tilde{\pi}}{\bm{\mathbf{{q}}}}{\bm{\mathbf{{q}}}}^{\sf T}]={\bm{\mathbf{{G}}}}_{R}-\mathbb{E}[\frac{1}{\tilde{\pi}}{\bm{\mathbf{{q}}}}{\bm{\mathbf{{q}}}}^{\sf T}]={\bm{\mathbf{{0}}}}_{k\times k}$
Second, bound $\|{\bm{\mathbf{{X}}}}\|_{2}\leq\gamma$. Observe $\frac{1}{\tilde{\pi}}{\bm{\mathbf{{q}}}}{\bm{\mathbf{{q}}}}^{\sf T}$ is a rank-1 matrix with its 2-norm bounded by $\max_{i\in{\cal I}_{R}}\frac{\|{\bm{\mathbf{{u}}}}_{i}\|_{2}^{2}}{\tilde{p}_{i}}=\xi$. Then
$\|{\bm{\mathbf{{X}}}}\|_{2}\leq\|{\bm{\mathbf{{I}}}}_{k}-({\bm{\mathbf{{S}}}}_{D}{\bm{\mathbf{{U}}}}_{D})^{\sf T}({\bm{\mathbf{{S}}}}_{D}{\bm{\mathbf{{U}}}}_{D})\|_{2}+\|\frac{1}{\tilde{\pi}}{\bm{\mathbf{{q}}}}{\bm{\mathbf{{q}}}}^{\sf T}\|_{2}\leq 1+\xi.$
@@
Finally, we show $\|\mathbb{E}[{\bm{\mathbf{{X}}}}^{\sf T}{\bm{\mathbf{{X}}}}]\|\leq\nu^{2}$. By expanding $\mathbb{E}[{\bm{\mathbf{{X}}}}^{\sf T}{\bm{\mathbf{{X}}}}]$ and using the fact that
@@
(31) $Pr[\|{\bm{\mathbf{{I}}}}-{\bm{\mathbf{{U}}}}_{A}^{\sf T}{\bm{\mathbf{{S}}}}_{H}^{\sf T}{\bm{\mathbf{{S}}}}_{H}{\bm{\mathbf{{U}}}}_{A}\|_{2}^{2}\geq\epsilon_{s}]=\delta\leq 2k\exp\Big{[}-s\epsilon_{s}^{2}/\Big{(}2(|\xi-1|)+2(1+\xi)\epsilon_{s}/3\Big{)}\Big{]}$
for some desired probability $\delta$ and desired error tolerance $\epsilon_{s}$. By choosing
$\frac{144\xi\log(2k/\delta)}{\epsilon_{s}^{2}}\leq s_{R}$ we have with probability at least $1-\delta$